40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1399     Entry time: Fri Mar 13 05:16:21 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Locking update 
Yoichi, Osamu,

With adjustments of the loop gains during the CARM offset reduction, the IFO reaches arm_power = 25 sort of robustly unless the 3.8kHz oscillation rings up.
At arm_power = 25, the CARM and DARM start to oscillate at around 400Hz. Probably I need more gain tweaks.
Annoying thing is that the 3.8kHz oscillation sometimes rings up suddenly and kills the lock.
This can happen anywhere above arm_power = 6 or so.
Because of a strange structure in the CARM loop gain around 3.8kHz, we cannot increase the CARM UGF beyond 1kHz.
The attached plots are the AO path open loop transfer function (attm2 is the zoom of attm1) measured at arm_power = 13.

Tomorrow, I will lock the X-arm and measure the transfer function from the AO path input to the X-arm error signal to see
if there is the same structure at 3.8kHz (X-arm error signal has the 3.8kHz peak).
Attachment 1: AOTF2.png  47 kB  | Hide | Hide all
AOTF2.png
Attachment 2: AOTF2-zoom.png  45 kB  | Hide | Hide all
AOTF2-zoom.png
ELOG V3.1.3-